The acceleration of a particle is given as (-2t+12t^2)m/s^2 find the distance covered by the particle between time=0s and time=20s
1
Expert's answer
2018-06-07T07:46:08-0400
v=-t^2+4t^3
d=t^4-t^3/3
d(20)=〖20〗^4-〖20〗^3/3=13.3 m.